∫ sec^n β tan β dβ


1
Expert's answer
2021-04-15T06:56:09-0400

"=\int \frac{\sin \left(β\right)}{\cos ^{n+1}\left(β\right)}dβ"

"=\int \frac{\sin \left(β\right)}{\cos ^{n+1}\left(β\right)}dβ"

"=\int \:-\frac{1}{u^{n+1}}du"

"\frac{1}{u^{n+1}}=u^{-n-1}"

"=-\int \:u^{-n-1}du"

"=-\frac{u^{-n-1+1}}{-n-1+1}"

"=-\frac{\cos ^{-n-1+1}\left(β\right)}{-n-1+1}"

"=\frac{1}{n}\cos ^{-n}\left(β\right)"

"=\frac{1}{n}\cos ^{-n}\left(β\right)+C"